OpenStudy (anonymous):

find d^2(y)/d(x^2) in terms of x and y (second derivative) 1-xy=x-y

OpenStudy (anonymous):

\[1(\frac{ dy }{ dx })-(x)\frac{ dy }{ dx}+\frac{ dx }{dx }y=\frac{ dx }{ dx }-\frac{ dy }{ dx }\] tell me how this looks it's been a while since I differentiated with this method.

OpenStudy (anonymous):

this would be the setup to solve for the first derivative

OpenStudy (anonymous):

y=-1 & y"=0

OpenStudy (anonymous):

http://i.imgur.com/CMgMz.gif

OpenStudy (anonymous):

mother of god...

OpenStudy (anonymous):

how did u get the second derivative? @mahmit2012

OpenStudy (anonymous):

-( x*y' + y) = 1 -y'

OpenStudy (anonymous):

solve for y'

OpenStudy (anonymous):

differentiate the expression again

OpenStudy (anonymous):

I didn't do the second derivative yet. just do the first then repeat.

OpenStudy (anonymous):

if you get another y' in your expression for y" sub.s in your result from the first differentiation...

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

y=-1 y"=0 why you did not pay attention!

OpenStudy (anonymous):

x#1

OpenStudy (anonymous):

i got y' is -y-1/x-1 then i tried doing quotient rule and thne i got confused

OpenStudy (anonymous):

you fail at math @mahmit2012

OpenStudy (anonymous):

k

OpenStudy (anonymous):

you got y' right

OpenStudy (anonymous):

every body sleeping !! cool

OpenStudy (anonymous):

yeah -y-1/x-1

OpenStudy (anonymous):

now, f = ( -y-1) g = (x-1) (f/g)' = f'g -fg'/g^2 f' = -y' g' = 1 g^2 = (x-1)^2

OpenStudy (anonymous):

plug them in and use the result from the first diff. : y' = -y-1/x-1 to sub.s in for y'

OpenStudy (anonymous):

find y=-1 it's so easy !!!

OpenStudy (anonymous):

got it @blackrose636 ?

OpenStudy (anonymous):

how did u even get y=-1?

OpenStudy (anonymous):

gl @Callisto

OpenStudy (anonymous):

y-xy=x-1 so y=-1 for all x#1

OpenStudy (anonymous):

@Algebraic! did you get it !

OpenStudy (callisto):

Ah... 1 - xy = x-y y-xy = x-1 y (1-x) = x-1 y = (x-1) / (1-x) y=-1!

OpenStudy (anonymous):

Thank you @Callisto

OpenStudy (anonymous):

this is different seeing !